OpenStudy (anonymous):

find the derivate of f(x)= square root(x-5) using the tangent definition.

OpenStudy (amistre64):

what is a tangent definition?

OpenStudy (amistre64):

if you mean limits ... use a conjugate

OpenStudy (anonymous):

yep limits

OpenStudy (anonymous):

with \[\lim_{x \rightarrow 0}\]

OpenStudy (amistre64):

\[\lim\frac{f(x+h)-f(x)}{h}\] \[\lim\frac{f(x+h)-f(x)}{h}\frac{f(x+h)+f(x)}{f(x+h)+f(x)}\] \[\lim\frac{1}{h(f(x+h)+f(x))}\] \[\lim\frac{1}{f(x+h)+f(x)}\] \[\frac{1}{f(x+0)+f(x)}\] \[\frac{1}{2f(x)}\]

OpenStudy (amistre64):

the nature of a sqrt allows that to happen

OpenStudy (amistre64):

got a typo after the conjugate :) should be an h to cancel, not a 1

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

thank u very much i got it now

OpenStudy (amistre64):

....just for clean up \[\lim\frac{f(x+h)-f(x)}{h}\frac{f(x+h)+f(x)}{f(x+h)+f(x)}\] \[\lim\frac{f^2(x+h)-f^2(x)}{h[f(x+h)+f(x)]}\] \[since~f(x) = \sqrt{(x-5)}\] \[\lim\frac{h}{h[f(x+h)+f(x)]}\] \[\lim\frac{1}{f(x+h)+f(x)}\] etc ...

OpenStudy (anonymous):

n the answer is wat i told u first 1/(power of function) *function with (power -1)*its deriviate pwer is 2 function is x-5 ndits deriviate is 1. so end result is 1/(2(x-5)^(1/2))